OpenStudy (anonymous):

Identify the value of the x-intercept by estimating -b/m...y=2x-2 The y intercept is (0, -2)

OpenStudy (anonymous):

x-intercept is 1. Just put y=0 and you get the answer!

OpenStudy (anonymous):

from...y=2x-2 slope=2, b=-2 x-intercept = -b/m=-(-2)/2=1 ans..